OpenStudy (anonymous):

Determine whether this series is convergent or divergent by expressing sn as a telescoping sum. {n=2(sum)infinity} 2/(n^2-1)

OpenStudy (zarkon):

\[\frac{2}{n^2-1}=\frac{1}{n-1}-\frac{1}{n+1}\]

OpenStudy (anonymous):

Awesome - thanks :)